# Risk-Calibrated Affective Speech Recognition via Conformal Coverage Guarantees: A Stochastic Calibrative Framework for Emergent Uncertainty Quantification

### v1

@@ Abstract
Traffic safety challenges arising from extreme driver emotions highlight the urgent need for reliable emotion recognition systems. Traditional deep learning approaches in speech emotion recognition suffer from overfitting and poorly calibrated confidence estimates. We propose a framework integrating Conformal Prediction (CP) and Risk Control, using Mel-spectrogram features processed through a pre-trained convolutional neural network. Our key innovation is the development of a nonconformity score that heuristically measures how closely a classifier’s predictions align with given inputs. Through calibration samples, we compute this score and derive a statistically rigorous threshold based on user-specified risk level $\alpha$, constructing prediction sets with provable coverage guarantees ($\geq 1-\alpha$). The Risk Control framework enables task-specific adaptation through customizable loss functions, dynamically adjusting prediction set sizes while maintaining coverage guarantees. Cross-dataset experiments on IEMOCAP and TESS demonstrate: 1) Strict coverage guarantee, 2) Significant negative correlation between Average Prediction Set Size (APSS) and $\alpha$, revealing reduced model uncertainty under high-risk conditions. We further propose APSS as a novel metric for evaluating classification uncertainty. This approach enhances speech emotion recognition reliability, with direct applications in intelligent transportation systems and real-time emotion monitoring.
keywords:
Speech emotion recognition; Conformal Prediction; Risk Control; Mel spectrum; Prediction set coverage guarantee
[…]
[label1] organization=School of Automation Science and Electrical Engineering, Beihang University, addressline=No. 37 Xueyuan Road, Haidian District, city=Beijing, postcode=100191, state=Haidian District, country=China
@@ 1 Introduction
With the rapid development of deep learning technology, speech emotion recognition tasks based on convolutional neural networks have made significant progress. This paper extracts the Mel-spectrogram features of speech signals to characterize different emotion categories and conducts experimental comparisons. Despite this, the model may still encounter overfitting problems during training, resulting in performance degradation on new and unseen data. In addition, current deep learning models generally lack confidence or uncertainty measures for prediction results, which means that in practical applications, the model’s predictions may be unreliable. Recent studies have pointed out that the miscalibration problem still exists. For example, Zhao et al. (2023) pointed out that although convolutional neural networks show high accuracy in speech emotion recognition tasks, the confidence scores of their prediction results are often too high and fail to truly reflect the actual performance of the model, resulting in incorrect emotion recognition. Therefore, improving the uncertainty measure of the model has become the key to improving the performance of speech emotion recognition systems Psaros et al. (2023); Wang et al. (2025b).
@@
Although Conformal Prediction can effectively provide coverage guarantee, it only solves the coverage problem of the prediction set and fails to consider the reliability of the model under different risk levels. In order to solve this problem, the Risk Control framework is further introduced. The core idea of this framework is to adjust the risk level of the model on the basis of maintaining coverage guarantee to cope with various uncertainties that may arise in practical applications. Under the Risk Control framework, the beta value is used to quickly determine the optimal prediction set size through the binary search method. By processing the calibration set, the size of the prediction set can be adjusted under the constraint of a given risk level to meet the needs of risk control. At the same time, the loss function is also improved so that it not only focuses on the prediction error rate, but also considers the size and uncertainty of each prediction set. The improvement of the loss function enables the framework to be flexibly applied to different tasks and scenarios, such as emotion recognition, intelligent traffic monitoring, etc. By combining Conformal Prediction and Risk Control, not only the coverage of the model is guaranteed, but also the robustness and accuracy of the model are improved, which can adapt to different practical application needs. Experimental results show that this framework can maintain high-precision emotion recognition at different risk levels while reducing the overfitting and uncertainty of the model, making it more valuable for practical applications.
@@
In order to verify the effectiveness of the framework, a large number of experiments were conducted. First, starting from the pre-experimental part, the performance of different feature selection methods and pre-trained models was tested. In the pre-experimental stage, a variety of feature extraction methods were compared, including mel-spectrogram features, chromaticity features, and spectral contrast features. Through testing on the TESS dataset, it was found that the performance of the mel-spectrogram feature was better than other feature extraction methods, reaching a test set accuracy of 99.82%. Therefore, the mel-spectrogram feature was finally selected as the speech feature extraction method. Next, several different neural network models, including ResNet50, MobileNetV3, ShuffleNetV2, etc., were trained on the IEMOCAP dataset, and their performance in emotion recognition tasks was tested. Although the accuracy of these models on the training set reached 100%, the accuracy on the test set decreased to 44.75%, 40.56% and 35.06% respectively. This phenomenon shows that the model has a certain overfitting problem when facing unseen data, so the Conformal Prediction framework was introduced to ensure the reliability and statistical significance of the model prediction results. To prove the coverage guarantee of the framework, two rounds of validation were conducted. The first round was conducted on the same dataset to ensure that the model can provide stable performance on the calibration set and the test set, verifying that Conformal Prediction can ensure the accuracy of the prediction results at a given risk level. The second round of experiments conducted cross-dataset validation, that is, the model trained on the IEMOCAP dataset was applied to the TESS dataset to test the generalization ability of the model. The experimental results show that the model can still maintain a high coverage across different datasets and can effectively adapt to different data distributions, verifying the robustness of the framework. Next, the uncertainty of different models is evaluated based on the prediction set size. During the training and testing process of each model, the prediction set size is calculated, and the uncertainty of the model is quantified based on the prediction set size at each alpha value. By conducting multiple experiments under approximately the same parameter conditions, when the uncertainty of the model increases, the prediction set size usually increases. This shows that a larger prediction set means that the model has lower confidence in the recognition of certain emotion categories. In order to control this uncertainty, the Risk Control framework is adopted, combined with an improved loss function to optimize the prediction set size to ensure that the model can provide reliable prediction results at different risk levels. Through these experiments, we not only verified the effectiveness of Conformal Prediction in speech emotion recognition, but also proved the application value of Risk Control framework in reducing uncertainty and improving prediction accuracy. The final experimental results show that the combination of these two frameworks can maintain the high accuracy and robustness of the model under different data sets and risk levels, ensuring that the emotion recognition system can meet the needs of practical applications.
@@ 2.2 Conformal Prediction
Conformal Prediction Angelopoulos and Bates (2021); Ye et al. (2024); Barber et al. (2023); Wang et al. (2024); Sun and Yu (2023); Wang et al. (2025a) is an uncertainty quantification method that provides strict statistical guarantees for machine learning models. Unlike traditional point prediction, CP provides a set of possible prediction labels for each input sample, rather than just a single prediction label, by generating a prediction set containing the prediction results. The core idea is to calculate the nonconformity score to measure the difference between each predicted label and the model’s prediction result, thereby deciding whether the label should be included in the prediction set. The nonconformity score measures the model’s confidence in the prediction result. For classification tasks, it is based on the difference between the model’s predicted probability of the true category and the actual category; for regression tasks, it is the difference between the predicted value and the true value. Before generating the prediction set, Conformal Prediction first uses a calibration set to estimate the distribution of the nonconformity score. The calibration set is a part of the data separated from the training set and is specifically used to measure the prediction uncertainty of the model. By calculating the nonconformity score in the calibration set, Conformal Prediction generates a quantitative value that defines which prediction labels the model should include at a given confidence level. In the test phase, for each new sample, the model outputs a predicted label and its corresponding nonconformity score. Then, Conformal Prediction generates a prediction set by comparing the inconsistency score of each label with the quantized value in the calibration set. The prediction set contains all labels whose inconsistency scores are less than or equal to the quantized value. In this way, the generated prediction set can ensure that the true label always appears in the prediction set at the specified confidence level. This provides a coverage guarantee, that is, at the set confidence level, the probability that the generated prediction set contains the true label is at least the confidence level (such as 90%). This guarantee applies to all models and does not depend on any assumed data distribution, so it is called a distribution-independent prediction method.
@@ 3.1 Preliminaries
As a statistical consistency prediction method, the SCP (Split Conformal Prediction) framework shows unique advantages in the reliability evaluation of machine learning models by converting heuristic uncertainty measures into prediction sets with strict probability guarantees. Its core idea is based on the assumption of data exchangeability and combines the quantile calibration of nonconformity scores to achieve coverage probability control of unknown samples. The theoretical basis and algorithm flow of SCP will be systematically explained below.
@@ 3.1.2 Mathematical Proof Supplement
3.2 Risk Control in speech emotion recognition task

### v2

# Coverage-Guaranteed Speech Emotion Recognition via Calibrated Uncertainty-Adaptive Prediction Sets

@@ Abstract
Road rage, stemming from emotional suppression and outbursts, can lead to traffic incidents such as rear-end collisions, crashes, and aggressive confrontations, posing significant risks to road safety and public security. Speech Emotion Recognition (SER) methods can detect early signs of negative emotions and provide real-time alerts to drivers, effectively reducing the likelihood of accidents. However, previous studies, such as Hidden Markov Models (HMMs) and Long Short-Term Memory (LSTM) networks, predominantly rely on one-dimensional speech signals, and are prone to overfitting and miscalibration issues, undermining their reliability in high-stakes applications. In this paper, we introduce a risk management framework that produces prediction sets with statistically rigorous guarantees of correctness coverage marginally over the test set. Specifically, we separate a held-out calibration set and develop a binary loss function for each data point to determine whether the ground-truth label is covered by the prediction set, which is calibrated by a data-driven threshold $\lambda$. We develop a joint loss function on the calibration set, and given a user-specified risk level $\alpha$, we then calibrate the threshold $\lambda$ to strictly bound it such that we can manage the expectation of the loss of each test point by this predefined upper bound $\alpha$. We conduct extensive evaluations on 6 base models on 2 datasets. Empirical results demonstrate that we can always bound the average correctness coverage rate (i.e., $\geq 1-\alpha$) at various user-specified risk levels. Moreover, we consistently control the marginal error rate across various splitting ratios (e.g., 0.1) between the calibration and test set, highlighting the robustness and effectiveness of our framework. Furthermore, we propose a small batch online calibration framework based on the local exchangeability assumption, which is suitable for complex scenarios where there is data domain offset or non-IID between batches. By constructing a non-negative test martingale, we ensure the coverage of the prediction set in dynamic scenarios, and verify the effectiveness of this method in controlling the coverage of the prediction set through cross-dataset experiments.
@@ 1 Introduction
With the rapid development of deep learning technology, speech emotion recognition tasks based on convolutional neural networks have made significant progress. This paper extracts the Mel-spectrogram features of speech signals to characterize different emotion categories and conducts experimental comparisons. Despite this, the model may still encounter overfitting problems during training, resulting in performance degradation on new and unseen data. In addition, current deep learning models generally lack confidence or uncertainty measures for prediction results, which means that in practical applications, the model’s predictions may be unreliable. Recent studies have pointed out that the miscalibration problem still exists. For example, Zhao et al. (2023) pointed out that although convolutional neural networks show high accuracy in speech emotion recognition tasks, the confidence scores of their prediction results are often too high and fail to truly reflect the actual performance of the model, resulting in incorrect emotion recognition. Therefore, improving the uncertainty measure of the model has become the key to improving the performance of speech emotion recognition systems Psaros et al. (2023); Wang et al. (2025a).
@@
Although CP can effectively provide coverage guarantee, it only solves the coverage problem of the prediction set without considering the reliability of the model under different risk levels. In order to solve this problem, a risk control framework is further introduced. The core idea of this framework is to adjust the risk level of the model on the basis of maintaining coverage guarantee to cope with various uncertainties that may arise in practical applications. Under the risk control framework, the beta value is used to quickly determine the optimal prediction set size through binary search. By processing the calibration set, the size of the prediction set can be adjusted under the constraint of a given risk level to meet the needs of risk control. At the same time, the loss function has been improved so that it not only focuses on the prediction error rate, but also considers the size of each prediction set and its uncertainty. The improvement of the loss function enables the framework to be flexibly applied to different tasks and scenarios, such as emotion recognition, intelligent traffic monitoring, etc. By combining CP with risk control, not only the coverage of the model is guaranteed, but also the robustness and accuracy of the model are improved, so that it can adapt to different practical application needs. Experimental results show that the framework can achieve coverage guarantee at different risk levels, while reducing the uncertainty of the model to a certain extent, making it more valuable for practical applications. However, traditional CP methods rely on a strict assumption of data exchangeability, which is valid in controlled environments but often fails in dynamic real-world scenarios, such as speech emotion recognition, where the data distribution may change between batches or domains. Such a non-exchangeable environment leads to a degradation of coverage guarantees, thus limiting the practical application of CP. To overcome this limitation, we propose a novel framework to ensure the coverage of the prediction set in dynamic scenarios by performing conformal prediction via a mini-batch online calibration mechanism. Specifically, we relax the global exchangeability assumption by adopting a local exchangeability assumption within a mini-batch. By constructing a non-negative test martingale, our method ensures coverage even when the distribution changes between batches, such as speaker differences or acoustic environment differences.
@@
In order to verify the effectiveness of the framework, a large number of experiments were carried out. First, starting from the pre-experimental part, the performance of different feature selection methods and pre-trained models was tested. In the pre-experimental stage, a variety of feature extraction methods were compared, including mel-spectrogram features, chromaticity features, and spectral contrast features. Through testing on the TESS dataset, it was found that the performance of mel-spectrogram features was better than other feature extraction methods, achieving a test set accuracy of 99.82%. Therefore, mel-spectrogram features were finally selected as the speech feature extraction method. Next, several different neural network models, including ResNet, MobileNetV3, ShuffleNetV2, etc., were trained on the IEMOCAP dataset, and their performance on emotion recognition tasks was tested. Although the accuracy of these models on the training set reached 100%, the accuracy on the test set dropped to 41.60%, 31.94%, and 34.13%, respectively. This phenomenon indicates that the model has a certain overfitting problem when facing unseen data, so the Conformal Prediction framework is introduced to ensure the reliability and statistical significance of the model prediction results. In order to prove the coverage guarantee of the framework, two rounds of verification were carried out. The first round of experiments was conducted on the same dataset to ensure that the model can provide stable performance on both the calibration set and the test set, verifying that Conformal Prediction can guarantee the accuracy of the prediction results under a given risk level. The second round of experiments conducted cross-dataset validation, that is, the model trained on the IEMOCAP dataset was applied to the TESS dataset to test the generalization ability of the model. The experimental results show that the model can still maintain a high coverage on different datasets and can effectively adapt to different data distributions, verifying the robustness of the framework. Next, the uncertainty of different models is evaluated based on the prediction set size. During the training and testing process of each model, the prediction set size is calculated, and the uncertainty of the model is quantified according to the prediction set size under each alpha value. By conducting multiple experiments under approximately the same parameter conditions, we found that when the uncertainty of the model increases, the prediction set size usually increases. This shows that a larger prediction set means that the model has lower confidence in the recognition of certain emotion categories. To control this uncertainty, we adopt the risk control framework to conduct the same experiment, and combine it with the improved loss function to optimize the prediction set size to ensure that the model can provide reliable prediction results under different risk levels. The final experimental results show that both frameworks can maintain the coverage guarantee and robustness of the final results under different datasets and risk levels, ensuring that the emotion recognition system can meet the needs of practical applications. Crucially, to verify our small batch martingale value method in the scenario of non-IID and domain shift, we designed an additional experiment in which the calibration and test data come from different datasets (IEMOCAP for calibration and TESS for testing), explicitly violating global exchangeability. The data is split into small batches (5 calibration samples + 1 test sample per batch), and the small amount of data within the batch is assumed to be locally exchangeable, while allowing for distribution shifts across batches. The coverage of traditional CP methods that rely on global exchangeability will drop significantly. In contrast, our martingale-based method brings the coverage close to the theoretical bound. This experiment finally proves the effectiveness of our framework in the scenario of non-IID and domain shift.
@@ 2.2 Conformal Prediction
CP Angelopoulos and Bates (2021); Ye et al. (2024); Barber et al. (2023); Wang et al. (2024); Sun and Yu (2023); Wang et al. (2025b) is an uncertainty quantification method that provides strict statistical guarantees for machine learning models. Unlike traditional point prediction, CP provides a set of possible prediction labels for each input sample, rather than just a single prediction label, by generating a prediction set containing the prediction results. The core idea is to calculate the nonconformity score to measure the difference between each predicted label and the model’s prediction result, thereby deciding whether the label should be included in the prediction set. The nonconformity score measures the model’s confidence in the prediction result. For classification tasks, it is based on the difference between the model’s predicted probability of the true category and the actual category; for regression tasks, it is the difference between the predicted value and the true value. Before generating the prediction set, Conformal Prediction first uses a calibration set to estimate the distribution of the nonconformity score. The calibration set is a part of the data separated from the training set and is specifically used to measure the prediction uncertainty of the model. By calculating the nonconformity score in the calibration set, Conformal Prediction generates a quantitative value that defines which prediction labels the model should include at a given confidence level. In the test phase, for each new sample, the model outputs a predicted label and its corresponding nonconformity score. Then, CP generates a prediction set by comparing the inconsistency score of each label with the quantized value in the calibration set. The prediction set contains all labels whose inconsistency scores are less than or equal to the quantized value. In this way, the generated prediction set can ensure that the true label always appears in the prediction set at the specified confidence level. This provides a coverage guarantee, that is, at the set confidence level, the probability that the generated prediction set contains the true label is at least the confidence level (such as 90%). This guarantee applies to all models and does not depend on any assumed data distribution, so it is called a distribution-independent prediction method.
@@ 3.1 Preliminaries
As a statistical consistency prediction method, the Split Conformal Prediction(SCP) framework shows unique advantages in the reliability evaluation of machine learning models by converting heuristic uncertainty measures into prediction sets with strict probability guarantees. Its core idea is based on the assumption of data exchangeability and combines the quantile calibration of nonconformity scores to achieve coverage probability control of unknown samples. The theoretical basis and algorithm flow of SCP will be systematically explained below.
@@ 3.1.2 Mathematical Proof Supplement
3.2 Risk Control in speech emotion recognition task1
@@ 3.3 Comparison of two frameworks
3.4 Mini-batch conformal prediction theory under non-IID1
In light of the limitations of traditional conformal prediction, which relies on global data exchangeability Barber et al. (2023); Kang et al. (2024); Wang et al. (2025c), we propose the assumption of within - batch local exchangeability. We prove that even when there are differences in data distributions between batches (i.e., batch - to - batch exchangeability may not hold), we can still strictly guarantee the global coverage probability by constructing a non - negative test supermartingale with a forcing operation. This framework breaks through the strong dependence of classical theories on independent and identically distributed data, providing universal theoretical support for small - batch scenarios such as time - series data and online learning.
E - value Gauthier et al. (2025), as a non - negative random variable satisfying $E[E]\leq 1$, can directly construct conformal sets through probability inequalities, offering a more flexible uncertainty quantification tool. For the calibration set $D_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ and the test sample $(X_{n+1},Y_{n+1})$, the E - value is defined as:
$E=\frac{S_{n+1}}{\frac{1}{n+1}\sum_{i=1}^{n+1}S_{i}}$ (12)
where $S_{i}=S(X_{i},Y_{i})$ and $S$ is a non - conformity score (such as cross - entropy loss) with non - negative values. Under the assumption of within - batch data exchangeability, the conditional expectation of the E - value satisfies $E[E_{t}|F_{t-1}]=1$, which provides a crucial basis for constructing the martingale process.
The filtration $\{F_{t}\}_{t\geq 0}$ is defined as the $\sigma$ algebra generated by all historical data up to the $t$ batch. Specifically, we define the filtration
$F_{t}=\sigma(S_{1}^{1},\cdots,S_{n_{1}}^{1},S_{n_{1}+1}^{1},\cdots,S_{1}^{t},\cdots,S_{n_{t}}^{t},S_{n_{t}+1}^{t})$ (13)
For all $t\geq 0$, the sequence of random variables $\{M_{t}\}_{t\geq 0}$ is defined as:
$M_{t}=\max\{(1-\gamma)E_{t}^{m_{-}true}+\gamma M_{t-1},1\}$ (14)
where
$E_{t}^{m_{-}true}=\frac{(n_{t}+1)\times s_{n_{t}+1}^{m_{-}true}}{\sum_{i=1}^{n_{t}}s_{i}+s_{n_{t}+1}^{m_{-}true}}$ (15)
We will prove that $\{M_{t}\}_{t\geq 0}$ is a non - negative test supermartingale.
  • Initial condition: First, we define $M_{0}=1$. Since the non - conformity scores we calculate are positive, it is obvious that for all $t\geq 0$, $M_{t}\geq 0$. The remaining task is to prove that $\{M_{t}\}_{t\geq 0}$ is a supermartingale. A non - negative test supermartingale needs to satisfy: $M_{0}=1$ (already satisfied), $M_{t}\geq 0$ and is $F_{t}$ measurable, and $E[M_{t}|F_{t-1}]\leq M_{t-1}$.
  • Measurability: To determine whether $M_{t}$ is measurable after defining the filtration, we use mathematical induction. When $t=0$, $M_{0}=1$. Since a constant function is measurable with respect to any $\sigma$ algebra, $M_{0}$ is $F_{0}$ measurable, where $F_{0}=\{\varnothing,\Omega\}$ is the trivial $\sigma$ algebra. Assume that $M_{t-1}$ is $F_{t-1}$ measurable. Because $F_{t-1}\subseteq F_{t}$, $M_{t-1}$ is also $F_{t}$ measurable. For the $t$ batch, $E_{t}^{m_{-}true}$ only depends on the non - conformity scores $S_{1}^{t},\cdots,S_{n_{t}}^{t},S_{n_{t}+1}^{t}$ of the $t$ th batch, and these scores are all in $F_{t}$, so $E_{t}^{m_{-}true}$ is $F_{t}$ measurable. According to the update rule, let $Z_{t}=(1-\gamma)E_{t}^{m_{-}true}+\gamma M_{t-1}$. Since $E_{t}^{m_{-}true}$ and $M_{t-1}$ are both $F_{t}$ measurable, the linear combination $Z_{t}$ is also $F_{t}$ measurable. Let $A_{t}=\{\omega\in\Omega:Z_{t}(\omega)<1\}$. Since $Z_{t}$ is $F_{t}$ measurable, $A_{t}\in F_{t}$. We can rewrite $M_{t}$ as $M_{t}=1\cdot I_{A_{t}}+Z_{t}\cdot I_{A_{t}^{c}}$, where $I_{A_{t}}$ is the indicator function of the set $A_{t}$. The indicator functions $I_{A_{t}}$ and $I_{A_{t}^{c}}$ are both $F_{t}$ measurable, and the products of the constant 1 and the measurable function $Z_{t}$ with the indicator functions are also $F_{t}$ measurable. Thus, their sum $M_{t}$ is $F_{t}$ measurable. In conclusion, by mathematical induction, we have proved that for the defined filtration $\{F_{t}\}_{t\geq 0}$, $M_{t}$ is $F_{t}$ measurable.
  • Decreasing conditional expectation: Under the assumption of within - batch data exchangeability, $E[E_{t}^{m_{-}true}|F_{t-1}]=1$. Then the conditional expectation of the supermartingale is
    $E[M_{t}|F_{t-1}]=(1-\lambda)E[E_{t}^{m_{-}true}|F_{t-1}]+\lambda M_{t-1}=(1-\lambda)+\lambda M_{t-1}$ (16)
    Since $M_{t-1}\geq 1$ and $\lambda\in(0,1)$, we have $E[M_{t}|F_{t-1}]\leq M_{t-1}$, which satisfies the property of a supermartingale.
Ville’s inequality states that for a non - negative supermartingale $M_{t}$, $P(\forall t\geq 0,M_{t}<\frac{1}{\alpha})\geq 1-\alpha$. The construction rule of the prediction set is to include the corresponding category in the prediction set when $M_{t}<\frac{1}{\alpha}$. This means that the probability that the true label falls into the prediction set is at least $1-\alpha$, thus theoretically guaranteeing the coverage rate of the prediction set. Let $x_{t}$ be the $t$ th test sample and $Y$ be the set of all possible categories. For the test sample $x_{t}$, its prediction set $\hat{C}(x_{t})$ can be expressed as:
$\hat{C}(x_{t})=\left\{c\in Y:\max\{(1-\gamma)\frac{(n_{t}+1)\cdot s(x_{t},c)}{\sum_{i=1}^{n_{t}}s(x_{t},y_{i})+s(x_{t},c)}+\gamma M_{t-1},1\}<\frac{1}{\alpha}\right\}$ (17)
Algorithm 1 Adaptive Non-Conformal Prediction with Supermartingale (ANCP-S)
1:
2:Pre-trained classifier $f:\mathcal{X}\to\Delta^{K-1}$
3:Significance level $\alpha\in(0,1)$
4:Mixing rate $\gamma\in(0,1)$
5:Initialize
6:Supermartingale $M_{0}\leftarrow 1$
7:procedure OnlinePrediction(Test stream $\{(x_{t},y_{t})\}_{t=1}^{T}$, $\mathcal{D}_{\text{cal}}$)
8:     for $t=1$ to $T$ do
9:         Sample $\mathcal{B}_{t}\subset\mathcal{D}_{\text{cal}}$ with $|\mathcal{B}_{t}|=5$
10:         $\bm{p}_{t}\leftarrow f(x_{t})$triangleright $\bm{p}_{t}\in\Delta^{K-1}$
11:         Compute $\{s_{i}\}_{i=1}^{n}$ where $s_{i}=1-\bm{p}_{z_{i}}(y_{i})\ \forall z_{i}\in\mathcal{B}_{t}$
12:         $C_{t}\leftarrow\emptyset$
13:         for $m=1$ to $K$ do
14:              $s^{m}_{t}\leftarrow 1-\bm{p}_{t}(m)$
15:              $E^{m}_{t}\leftarrow\frac{(n+1)s^{m}_{t}}{\sum_{s_{i}\in\mathcal{B}_{t}}s_{i}+s^{m}_{t}}$
[…]
17:              if $M^{m}_{t}<1/\alpha$ then
18:                  $C_{t}\leftarrow C_{t}\cup\{m\}$
19:              end if
20:         end for
21:         $E^{y_{t}}_{t}\leftarrow\frac{(n+1)(1-\bm{p}_{t}(y_{t}))}{\sum_{s_{i}\in\mathcal{B}_{t}}s_{i}+(1-\bm{p}_{t}(y_{t}))}$
[…]
23:     end for
24:end procedure